# From Word Vectors to Multimodal Embeddings: Techniques, Applications, and Future Directions For Large Language Models

### v1

@@ Abstract
Index Terms:
Large Language Model, Word Embeddings, Contextualized Embeddings, Multimodal Representations, Natural Language Processing
@@ I Introduction
Large Language Models (LLMs) have transformed natural language processing (NLP) by providing advanced tools for understanding and generating human language. At the core of these models are word embeddings—dense, continuous vector representations that capture semantic and syntactic relationships among words. By mapping words into high-dimensional spaces where semantically related words are situated near each other, embeddings support nuanced language interpretation and have become essential to NLP applications such as machine translation, sentiment analysis, and information retrieval. The progression from early one-hot encodings to more sophisticated embeddings like Word2Vec, GloVe, and fastText has markedly improved the accuracy and scalability of language models, enabling them to handle large volumes of textual data with high precision [1, 2, 3].
Architectures like ELMo, BERT, and GPT employ deep neural networks to generate embeddings that reflect context-dependent meanings, addressing complexities such as polysemy and capturing long-range dependencies in language. These contextual representations not only enhance the accuracy and robustness of NLP systems but also support the integration of language with other modalities, including vision and robotics, enabling more sophisticated, interactive applications. However, despite these advancements, challenges persist concerning efficiency, interpretability, and ethical concerns within embedding models. High computational costs, the opaque nature of embedding spaces, and the risk of propagating biases from training data highlight areas for continued research [4, 5, 6]. This review examines the evolution of word embeddings, exploring foundational principles, varied methodologies, cross-modal applications, and key challenges.
{forest}
[…]
Concepts, level1, [Distributional Hypothesis [7, 8, 9], level2] [Contextual Similarity [10, 11, 12], level2] ] [From Sparse
to Dense
Representations, level1, [One-Hot Encoding [13, 10, 14], level2] [Word Embeddings [10, 8, 15], level2] ] [Contextualized
Word Embeddings, level1, [ELMo [16], level2] [BERT and Variants [17, 18, 19], level2] [Other Contextualized Embeddings:
GPT, XLNet, XLM [20, 21, 22], level2] ] [Subword-Level
Embeddings and
Generalization, level1, [Handling Rare and Unseen
Words [23, 24, 25, 10, 26], level2] [Cross-Lingual Embeddings [27, 28, 29, 30], level2] ] [Personalized
Word Embeddings, level1, [Modeling Linguistic Variation [31, 32], level2] [Applications [31, 11, 33], level2] ] ]
@@ II-A1 Distributional Hypothesis
The distributional hypothesis, a cornerstone of numerous word embedding techniques, posits that words appearing in similar contexts tend to have similar meanings [7]. This hypothesis allows for representing words as vectors in a continuous space, where semantic similarity is reflected by vector proximity [8]. This shift from symbolic to distributed representations has revolutionized NLP, enabling advancements in tasks like information retrieval, machine translation, and sentiment analysis [9]. The distributional hypothesis has limitations, however. It often struggles to capture the nuances of word meaning in different contexts, particularly for polysemous words, leading to a need for context-dependent representations [34]. Additionally, while effective at capturing broad semantic and syntactic relations, traditional word embeddings based on the distributional hypothesis can be computationally expensive and struggle with issues like the curse of dimensionality, out-of-vocabulary words, and overfitting to specific domains [15]. Moreover, focusing primarily on word properties rather than morphology can lead to inconsistent performance across different evaluation metrics [13]. This has motivated research into alternative word embedding models that incorporate subword information, particularly for morphologically rich languages [27].
@@ II-A2 Contextual Similarity
Context plays a critical role in disambiguating word meanings and improving the performance of language models. Different definitions of context offer unique approaches to capture semantic relationships. Local context, often defined as a sliding window of neighboring words, is utilized by models like Word2Vec to learn word embeddings through the prediction of nearby words [10]. Although efficient, this approach can be limited in its ability to capture long-range dependencies. Sentence-level context, which considers the entire sentence, allows models like LSTM-RNNs to integrate more comprehensive information, resulting in more nuanced sentence embeddings [11]. Document-level context further expands the scope, encompassing the entire document, which proves advantageous for tasks such as document classification and in representing temporal relationships between documents in sequential data streams [12].
Contextualized word embeddings, generated by models such as ELMo and BERT, incorporate contextual information directly into word representations and enable the same word to have distinct embeddings based on its context [35]. This contextualization is crucial for addressing polysemy and enhancing performance in tasks like word sense disambiguation (WSD) [16]. For instance, contextual string embeddings, which represent words as character sequences and integrate surrounding text, have proven especially effective for WSD, notably in named entity recognition tasks [36]. Additionally, the type and extent of context used can substantially influence language model performance. For example, while adding socio-situational information offers slight improvements to character-based models, it may adversely affect models using word embeddings, underscoring the need to carefully assess the interplay between context and embedding type [37]. This has driven research into refined context selection strategies and data augmentation techniques aimed at further enhancing language model capabilities [38].
@@ II-B1 One-Hot Encoding
One-hot encoding represents words as sparse, high-dimensional vectors. Each word in the vocabulary is assigned a unique vector in which only one element is set to one, while all others remain zero. Despite its simplicity, this representation has several critical limitations. The dimensionality of these vectors scales linearly with the vocabulary size, leading to extremely high dimensions for large vocabularies [13]. This results in significant sparsity, with most elements being zero, making computations inefficient and resource-intensive [10]. Additionally, one-hot encoding lacks the ability to capture semantic relationships between words. Each word is represented independently, with no intrinsic way to reflect similarities or relationships between words. This is a substantial limitation because capturing semantic and syntactic relationships is crucial for many natural language processing (NLP) tasks [39]. For instance, as discussed in [14], in an n-gram language model, predicting the next word in a sequence relies on word co-occurrence. If certain words never co-occur, the model cannot infer any relationship between them. Thus, one-hot encoding fails to capture the contextual similarity needed for effective language modeling and other NLP tasks, such as semantic search, knowledge base question answering, and machine translation [23, 40].
@@ II-B2 Word Embeddings
Word embeddings address the limitations of one-hot encoding by representing words as dense, low-dimensional vectors learned from large corpora [10]. These vectors capture semantic and syntactic relationships, mapping words into a continuous vector space where similar words are positioned closer together [8]. This arrangement is consistent with the distributional hypothesis, which states that words occurring in similar contexts tend to have similar meanings [41]. This contextual information is valuable for a range of NLP tasks, including word similarity measurement [13], analogy solving [42], document retrieval [11], and machine translation [30].
The continuous vector space representation also allows algebraic operations on word embeddings. This property is particularly useful in tasks such as analogy solving, where relationships between words can be expressed as vector operations [43]. By embedding words into a latent space, these models capture the underlying semantic and syntactic structure of language [15]. This latent representation is especially advantageous for downstream tasks like question answering, named entity recognition, and neural machine translation, as it enables models to generalize beyond specific word occurrences and recognize broader semantic relationships. Unlike one-hot encoding, word embeddings mitigate the problems of high dimensionality, sparsity, and the lack of semantic representation [39, 34]. The dense vectors allow more efficient computations and enable the model to learn complex relationships based on the distributional properties of large text corpora [26].
Word embeddings extend beyond individual words to represent phrases and even entire documents, as shown in [44] and [45]. This versatility makes word embeddings a powerful and flexible tool for a wide range of NLP applications.
@@ II-C1 ELMo
ELMo [16] employs bidirectional LSTMs trained on a language modeling objective to generate contextualized word representations. Unlike static word embeddings (e.g., Word2Vec, GloVe) that assign a single vector per word regardless of context, ELMo produces dynamic embeddings that vary based on the word’s surrounding text. This contextual sensitivity allows ELMo to capture nuanced meanings and disambiguate polysemous words. The architecture consists of two LSTM layers stacked on top of each other, each processing the input sequence in both forward and backward directions. The internal states of these LSTMs at each time step are combined, often through a weighted average, to form the contextualized word representation. Different layers of the bidirectional LSTM capture different aspects of linguistic information. The lower layers tend to encode syntactic information, while higher layers focus on semantic aspects. Alternative weighting schemes for combining these layers can be learned during downstream task training, optimizing the contribution of each layer for specific tasks.
@@ II-C2 BERT and Variants
BERT [17] and other transformer-based bidirectional encoders use a Transformer encoder architecture and are pre-trained on two objectives: Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). MLM randomly masks tokens in the input sequence and trains the model to predict the masked words based on the surrounding context. NSP trains the model to determine whether two given sentences are consecutive in the original text. These objectives enable BERT to capture bidirectional context effectively, addressing limitations of previous models that relied on unidirectional or shallow bidirectional representations. Fine-tuning BERT for specific downstream tasks involves adding task-specific layers on top of the pre-trained encoder and training the entire model on labeled data for the target task.
RoBERTa [18] modifies the BERT pre-training procedure by removing NSP, training with larger batch sizes and more data, and dynamically changing the masking pattern during training. ALBERT [19] introduces parameter reduction techniques, such as factorized embedding parameterization and cross-layer parameter sharing, to reduce the model size and improve training efficiency while maintaining performance. These variants address some of the computational challenges associated with BERT and offer improved performance on various downstream tasks. BERT and its variants handle polysemy by generating distinct embeddings for a word based on its context. The MLM objective helps capture the meaning of rare words by leveraging the surrounding context. However, long sequences can still pose challenges due to the quadratic complexity of the self-attention mechanism in the Transformer architecture. The tying of word vectors and word classifiers as a loss framework for language modeling [46] has also been explored as a way to improve language modeling performance.
@@ II-C3 Other Contextualized Embeddings: GPT, XLNet, and XLM
GPT [20] utilizes a Transformer decoder architecture and is trained using a language modeling objective, predicting the next word in a sequence. This autoregressive approach captures dependencies in one direction, limiting its ability to capture full bidirectional context. XLNet [21] addresses this limitation by using a permutation language modeling objective, which considers all possible orderings of the input sequence during training, allowing it to capture bidirectional context while maintaining the autoregressive formulation. XLM [22] extends BERT to support cross-lingual training, utilizing a translation language modeling objective to learn representations that capture relationships between words in different languages. The input embeddings of these models typically represent word-level information, while the output embeddings encode contextualized representations. Tying input and output embeddings [46] can be used to reduce the number of parameters and potentially improve performance. Using the output embedding to improve language models has been explored in various contexts, often by incorporating it into downstream tasks or as a way to refine the language model itself.
@@ II-D1 Subword Information for Handling Rare and Unseen Words
[23] examined word embeddings to expand queries with semantically related terms, enhancing the accuracy and robustness of query language models in handling vocabulary mismatches. Building on this, [24] proposed a generalized language model utilizing word embeddings to address vocabulary mismatch in information retrieval. This model conceptualizes query term observation as a two-step process: generating an intermediate term from the document or collection and transforming it into the observed query term, thus effectively capturing term relationships and mitigating vocabulary mismatches.
In a similar vein, [25] introduced a subword-level vector model that treats words as bags of character n-grams, allowing it to generate effective embeddings for rare or unseen words. This model is computationally efficient, easy to train, and demonstrates state-of-the-art performance on word similarity and morphosyntactic tasks across multiple languages. Complementing this approach, [10] presented a simplified method for training word embeddings on rare or unseen words, using noise-contrastive estimation for log-bilinear models. This method is faster and more efficient than prior models, producing better results in terms of embedding quality and generalization. Additionally, [26] offers an extensive review of word embedding models, including subword-level approaches, and their applications across various NLP tasks.
@@ II-D2 Cross-Lingual Word Embeddings and Low-Resource Languages
Subword information is instrumental in advancing cross-lingual word embeddings, particularly for low-resource languages with limited training data. By capturing morphological similarities across languages, subword-level models can learn shared representations for morphologically related words, even when they exhibit different surface forms. [27] explored universal and language-specific properties within word embeddings, revealing that word form features are particularly beneficial for inflectional languages. Similarly, [28] examined cross-lingual word embeddings derived from bilingual lexicons to enhance language models for low-resource languages. Applying this approach to Yongning Na, they highlighted challenges and potential solutions for low-resource settings.
Furthermore, [29] developed a language-agnostic BERT model, LaBSE, which supports over 100 languages and sets a new benchmark for cross-lingual tasks. LaBSE leverages both multilingual and monolingual data, using techniques like MLM and TLM to learn effective cross-lingual sentence embeddings. Finally, [30] introduced bilingual word embeddings based on a large, unlabeled corpus and machine translation word alignments, showing notable improvements in machine translation by capturing semantic equivalence across languages.

### v2

@@ I Introduction
Large Language Models (LLMs) have transformed natural language processing (NLP) by providing advanced tools for understanding and generating human language. At the core of these models are word embeddings—dense, continuous vector representations that capture semantic and syntactic relationships among words. By mapping words into high-dimensional spaces where semantically related words are situated near each other, embeddings support nuanced language interpretation and have become essential to NLP applications such as machine translation, sentiment analysis, and information retrieval. The progression from early one-hot encodings to more sophisticated embeddings like Word2Vec, GloVe, and fastText has markedly improved the accuracy and scalability of language models, enabling them to handle large volumes of textual data with high precision [rifat_robust_2020, victor_language_2016, novotn_y_when_2022].
Architectures like ELMo, BERT, and GPT employ deep neural networks to generate embeddings that reflect context-dependent meanings, addressing complexities such as polysemy and capturing long-range dependencies in language. These contextual representations not only enhance the accuracy and robustness of NLP systems but also support the integration of language with other modalities, including vision and robotics, enabling more sophisticated, interactive applications. However, despite these advancements, challenges persist concerning efficiency, interpretability, and ethical concerns within embedding models. High computational costs, the opaque nature of embedding spaces, and the risk of propagating biases from training data highlight areas for continued research [peng2024jailbreaking, peng2024securing, li2024surveying]. This review examines the evolution of word embeddings, exploring foundational principles, varied methodologies, cross-modal applications, and key challenges.
Word Embeddings Foundational Concepts Distributional Hypothesis [lebret_word_2016, xiaolei_low-dimensional_2019, sandip_comparative_2018] Contextual Similarity [a_learning_2013, hamid_deep_2015, nemanja_hierarchical_2015] From Sparse to Dense Representations One-Hot Encoding [yong_intrinsic_2018, a_learning_2013, rachith_inexplicable_2023] Word Embeddings [a_learning_2013, xiaolei_low-dimensional_2019, matthias_context_2018] Contextualized Word Embeddings ELMo [matthew_deep_2018] BERT and Variants [devlin2018bert, liu2019roberta, lan2019albert] Other Contextualized Embeddings: GPT, XLNet, XLM [radford2018improving, yang2019xlnet, conneau2019cross] Subword-Level Embeddings and Generalization Handling Rare and Unseen Words [hamed_embedding-based_2016, debasis_word_2015, jinman_generalizing_2018, a_learning_2013, mourad_word_2022] Cross-Lingual Embeddings [peng_investigating_2016, trevor_cross-lingual_2017, fangxiaoyu_language-agnostic_2020, will_bilingual_2013] Personalized Word Embeddings Modeling Linguistic Variation [charles_exploring_2020, yossi_analysis_2017] Applications [charles_exploring_2020, hamid_deep_2015, doddapaneni_user_2024]
@@ II-A1 Distributional Hypothesis
The distributional hypothesis, a cornerstone of numerous word embedding techniques, posits that words appearing in similar contexts tend to have similar meanings [lebret_word_2016]. This hypothesis allows for representing words as vectors in a continuous space, where semantic similarity is reflected by vector proximity [xiaolei_low-dimensional_2019]. This shift from symbolic to distributed representations has revolutionized NLP, enabling advancements in tasks like information retrieval, machine translation, and sentiment analysis [sandip_comparative_2018]. The distributional hypothesis has limitations, however. It often struggles to capture the nuances of word meaning in different contexts, particularly for polysemous words, leading to a need for context-dependent representations [tammie_introduction_2021]. Additionally, while effective at capturing broad semantic and syntactic relations, traditional word embeddings based on the distributional hypothesis can be computationally expensive and struggle with issues like the curse of dimensionality, out-of-vocabulary words, and overfitting to specific domains [matthias_context_2018]. Moreover, focusing primarily on word properties rather than morphology can lead to inconsistent performance across different evaluation metrics [yong_intrinsic_2018]. This has motivated research into alternative word embedding models that incorporate subword information, particularly for morphologically rich languages [peng_investigating_2016].
@@ II-A2 Contextual Similarity
Context plays a critical role in disambiguating word meanings and improving the performance of language models. Different definitions of context offer unique approaches to capture semantic relationships. Local context, often defined as a sliding window of neighboring words, is utilized by models like Word2Vec to learn word embeddings through the prediction of nearby words [a_learning_2013]. Although efficient, this approach can be limited in its ability to capture long-range dependencies. Sentence-level context, which considers the entire sentence, allows models like LSTM-RNNs to integrate more comprehensive information, resulting in more nuanced sentence embeddings [hamid_deep_2015]. Document-level context further expands the scope, encompassing the entire document, which proves advantageous for tasks such as document classification and in representing temporal relationships between documents in sequential data streams [nemanja_hierarchical_2015].
Contextualized word embeddings, generated by models such as ELMo and BERT, incorporate contextual information directly into word representations and enable the same word to have distinct embeddings based on its context [shao-yen_multimodal_2019]. This contextualization is crucial for addressing polysemy and enhancing performance in tasks like word sense disambiguation (WSD) [matthew_deep_2018]. For instance, contextual string embeddings, which represent words as character sequences and integrate surrounding text, have proven especially effective for WSD, notably in named entity recognition tasks [a_contextual_2018]. Additionally, the type and extent of context used can substantially influence language model performance. For example, while adding socio-situational information offers slight improvements to character-based models, it may adversely affect models using word embeddings, underscoring the need to carefully assess the interplay between context and embedding type [linda_effect_2019]. This has driven research into refined context selection strategies and data augmentation techniques aimed at further enhancing language model capabilities [matthias_scope_2018].
@@ II-B1 One-Hot Encoding
One-hot encoding represents words as sparse, high-dimensional vectors. Each word in the vocabulary is assigned a unique vector in which only one element is set to one, while all others remain zero. Despite its simplicity, this representation has several critical limitations. The dimensionality of these vectors scales linearly with the vocabulary size, leading to extremely high dimensions for large vocabularies [yong_intrinsic_2018]. This results in significant sparsity, with most elements being zero, making computations inefficient and resource-intensive [a_learning_2013]. Additionally, one-hot encoding lacks the ability to capture semantic relationships between words. Each word is represented independently, with no intrinsic way to reflect similarities or relationships between words. This is a substantial limitation because capturing semantic and syntactic relationships is crucial for many natural language processing (NLP) tasks [yang_word_2018]. For instance, as discussed in [rachith_inexplicable_2023], in an n-gram language model, predicting the next word in a sequence relies on word co-occurrence. If certain words never co-occur, the model cannot infer any relationship between them. Thus, one-hot encoding fails to capture the contextual similarity needed for effective language modeling and other NLP tasks, such as semantic search, knowledge base question answering, and machine translation [hamed_embedding-based_2016, japa_question_2020].
@@ II-B2 Word Embeddings
Word embeddings address the limitations of one-hot encoding by representing words as dense, low-dimensional vectors learned from large corpora [a_learning_2013]. These vectors capture semantic and syntactic relationships, mapping words into a continuous vector space where similar words are positioned closer together [xiaolei_low-dimensional_2019]. This arrangement is consistent with the distributional hypothesis, which states that words occurring in similar contexts tend to have similar meanings [marco_dont_2014]. This contextual information is valuable for a range of NLP tasks, including word similarity measurement [yong_intrinsic_2018], analogy solving [andrew_modeling_2015], document retrieval [hamid_deep_2015], and machine translation [will_bilingual_2013].
The continuous vector space representation also allows algebraic operations on word embeddings. This property is particularly useful in tasks such as analogy solving, where relationships between words can be expressed as vector operations [sanjeev_latent_2015]. By embedding words into a latent space, these models capture the underlying semantic and syntactic structure of language [matthias_context_2018]. This latent representation is especially advantageous for downstream tasks like question answering, named entity recognition, and neural machine translation, as it enables models to generalize beyond specific word occurrences and recognize broader semantic relationships. Unlike one-hot encoding, word embeddings mitigate the problems of high dimensionality, sparsity, and the lack of semantic representation [yang_word_2018, tammie_introduction_2021]. The dense vectors allow more efficient computations and enable the model to learn complex relationships based on the distributional properties of large text corpora [mourad_word_2022].
Word embeddings extend beyond individual words to represent phrases and even entire documents, as shown in [shaohua_embedding_2016] and [mohammad_embeddings_2021]. This versatility makes word embeddings a powerful and flexible tool for a wide range of NLP applications.
@@ II-C1 ELMo
ELMo [matthew_deep_2018] employs bidirectional LSTMs trained on a language modeling objective to generate contextualized word representations. Unlike static word embeddings (e.g., Word2Vec, GloVe) that assign a single vector per word regardless of context, ELMo produces dynamic embeddings that vary based on the word’s surrounding text. This contextual sensitivity allows ELMo to capture nuanced meanings and disambiguate polysemous words. The architecture consists of two LSTM layers stacked on top of each other, each processing the input sequence in both forward and backward directions. The internal states of these LSTMs at each time step are combined, often through a weighted average, to form the contextualized word representation. Different layers of the bidirectional LSTM capture different aspects of linguistic information. The lower layers tend to encode syntactic information, while higher layers focus on semantic aspects. Alternative weighting schemes for combining these layers can be learned during downstream task training, optimizing the contribution of each layer for specific tasks.
@@ II-C2 BERT and Variants
BERT [devlin2018bert] and other transformer-based bidirectional encoders use a Transformer encoder architecture and are pre-trained on two objectives: Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). MLM randomly masks tokens in the input sequence and trains the model to predict the masked words based on the surrounding context. NSP trains the model to determine whether two given sentences are consecutive in the original text. These objectives enable BERT to capture bidirectional context effectively, addressing limitations of previous models that relied on unidirectional or shallow bidirectional representations. Fine-tuning BERT for specific downstream tasks involves adding task-specific layers on top of the pre-trained encoder and training the entire model on labeled data for the target task.
RoBERTa [liu2019roberta] modifies the BERT pre-training procedure by removing NSP, training with larger batch sizes and more data, and dynamically changing the masking pattern during training. ALBERT [lan2019albert] introduces parameter reduction techniques, such as factorized embedding parameterization and cross-layer parameter sharing, to reduce the model size and improve training efficiency while maintaining performance. These variants address some of the computational challenges associated with BERT and offer improved performance on various downstream tasks. BERT and its variants handle polysemy by generating distinct embeddings for a word based on its context. The MLM objective helps capture the meaning of rare words by leveraging the surrounding context. However, long sequences can still pose challenges due to the quadratic complexity of the self-attention mechanism in the Transformer architecture. The tying of word vectors and word classifiers as a loss framework for language modeling [hakan_tying_2016] has also been explored as a way to improve language modeling performance.
@@ II-C3 Other Contextualized Embeddings: GPT, XLNet, and XLM
GPT [radford2018improving] utilizes a Transformer decoder architecture and is trained using a language modeling objective, predicting the next word in a sequence. This autoregressive approach captures dependencies in one direction, limiting its ability to capture full bidirectional context. XLNet [yang2019xlnet] addresses this limitation by using a permutation language modeling objective, which considers all possible orderings of the input sequence during training, allowing it to capture bidirectional context while maintaining the autoregressive formulation. XLM [conneau2019cross] extends BERT to support cross-lingual training, utilizing a translation language modeling objective to learn representations that capture relationships between words in different languages. The input embeddings of these models typically represent word-level information, while the output embeddings encode contextualized representations. Tying input and output embeddings [hakan_tying_2016] can be used to reduce the number of parameters and potentially improve performance. Using the output embedding to improve language models has been explored in various contexts, often by incorporating it into downstream tasks or as a way to refine the language model itself.
@@ II-D1 Subword Information for Handling Rare and Unseen Words
[hamed_embedding-based_2016] examined word embeddings to expand queries with semantically related terms, enhancing the accuracy and robustness of query language models in handling vocabulary mismatches. Building on this, [debasis_word_2015] proposed a generalized language model utilizing word embeddings to address vocabulary mismatch in information retrieval. This model conceptualizes query term observation as a two-step process: generating an intermediate term from the document or collection and transforming it into the observed query term, thus effectively capturing term relationships and mitigating vocabulary mismatches.
In a similar vein, [jinman_generalizing_2018] introduced a subword-level vector model that treats words as bags of character n-grams, allowing it to generate effective embeddings for rare or unseen words. This model is computationally efficient, easy to train, and demonstrates state-of-the-art performance on word similarity and morphosyntactic tasks across multiple languages. Complementing this approach, [a_learning_2013] presented a simplified method for training word embeddings on rare or unseen words, using noise-contrastive estimation for log-bilinear models. This method is faster and more efficient than prior models, producing better results in terms of embedding quality and generalization. Additionally, [mourad_word_2022] offers an extensive review of word embedding models, including subword-level approaches, and their applications across various NLP tasks.
@@ II-D2 Cross-Lingual Word Embeddings and Low-Resource Languages
Subword information is instrumental in advancing cross-lingual word embeddings, particularly for low-resource languages with limited training data. By capturing morphological similarities across languages, subword-level models can learn shared representations for morphologically related words, even when they exhibit different surface forms. [peng_investigating_2016] explored universal and language-specific properties within word embeddings, revealing that word form features are particularly beneficial for inflectional languages. Similarly, [trevor_cross-lingual_2017] examined cross-lingual word embeddings derived from bilingual lexicons to enhance language models for low-resource languages. Applying this approach to Yongning Na, they highlighted challenges and potential solutions for low-resource settings.
Furthermore, [fangxiaoyu_language-agnostic_2020] developed a language-agnostic BERT model, LaBSE, which supports over 100 languages and sets a new benchmark for cross-lingual tasks. LaBSE leverages both multilingual and monolingual data, using techniques like MLM and TLM to learn effective cross-lingual sentence embeddings. Finally, [will_bilingual_2013] introduced bilingual word embeddings based on a large, unlabeled corpus and machine translation word alignments, showing notable improvements in machine translation by capturing semantic equivalence across languages.